At the Horton College of Engineering, Engineering students and Physics students took a math exam. Their scores were used to make side-by-side boxplots. Answer the following questions: 1. Engineering o Physics 20 30 40 50 60 70 80 90 100 (1) What is the interquartile range among the physics students? (2) How low of a score on the exam would an engineering student need in order to be considered as an outlier? (3) Based on the boxplots, what is the distribution shape for engineering students and physics students scores, respectively? (i.e. right-skewed, symmetric, or left-screwed?)
Answer #1

1) interquartile range IQR=Q3-Q1=80-50=30

2)

here IQR for engineering student =90-70=20

hence lower outlier threshold =Q1-1.5*IQR=70-1.5*20=40 ; hence below 40 values are considered to be outlier

3)

for engineering student: symmetric

for physics student: left skewed